OpenStudy (anonymous):

(1 + √3)/(1 - √3) = (1 + √3)/(1 - √3) = @Mathematics

OpenStudy (lukecrayonz):

Equals math. answer, -2-sqrt(-3)

OpenStudy (anonymous):

can you show me how you got that please

OpenStudy (lukecrayonz):

sorry, -2-sqrt(3), NOT NEGATIVE 3.

OpenStudy (lukecrayonz):

Yeah saifoo howd you get -2 o.O

OpenStudy (anonymous):

luke can you show me how you got that

OpenStudy (anonymous):

here:

OpenStudy (saifoo.khan):

Oh wait, i thought that's multiplying! :O

OpenStudy (anonymous):

in order to get rid of the radical on the bottom, multiply both top and bottom by \[1+\sqrt{3}\]you will get\[(1+\sqrt{3})^{2} \over (1-\sqrt{3})(1+\sqrt{3})\]when you simplify you will get\[1+2\sqrt{3}+3 \over 1-3\]so\[4+2\sqrt{3} \over -2\]final answer:\[-2-\sqrt{3}\]

OpenStudy (anonymous):

thanks so much lia